A record of travel along a straight path is as follows:

1. Start from rest with constant acceleration of 2.40 m/s2 for 16.0 s.

2. Maintain a constant velocity for the next 1.20 min.

3. Apply a constant negative acceleration of −9.29 m/s2 for 4.13 s.

(a) What was the total displacement for the trip?

 



(b) What were the average speeds for legs 1, 2, and 3 of the trip, as well as for the complete trip?leg 1     



 m/sleg 3     


 m/scomplete trip     

 m/s


1
Expert's answer
2021-09-15T17:42:48-0400

a)

"l_1=\frac{a_1t_1^2}2=307.2~m,"

"v_1=a_1t_1=38.4~\frac ms,"

"l_2=v_1t_2=2764.8~m,"

"l_3=v_1t_3-\frac{a_2t_3^2}2=79.4~m,"

"l=l_1+l_3+l_3=3151.4~m,"

b)

"v_{av1}=\frac{v_1}2=19.2~\frac ms,"

"v_{av2}=v_1=38.4~\frac ms,"

"v_2=v_1-a_2t_3=0~\frac ms,"

"v_{av3}=\frac {v_1}2=19.2~\frac ms,"

"v_{av}=\frac lt=34.2~\frac ms."